1.      The equation in polar coordinates of an ellipse with semi-axes a and b is:





1
Expert's answer
2021-06-14T08:18:07-0400
"\frac{x^2}{a^2}+\frac{y^2}{b^2}=1"

"x=\rho cos\phi"

"y=\rho sin\phi"

"\rho=± \sqrt {\frac{a^2 b^2}{b^2cos^2\phi+a^2 sin^2 \phi}}"